OpenStudy (anonymous):

Solve the equation: x2 - 2x + 1 = 3.

jhonyy9 (jhonyy9):

do you know how you can solveing a quadratic equation ?

jhonyy9 (jhonyy9):

so first add to both sides -3 and will get x^2 -2x +1 -3 =3-3 x^2 -2x -2 =0 and now you need to solve it ,do you know it how ?